# Turning Waste into Wealth: Leveraging Low-Quality Samples for Enhancing Continuous Conditional Generative Adversarial Networks

### 14 Dec 2023 v2

@@ C.1 Some extra ablation studies
Ablation Study 1: The effect of various generators for generating Type II negative samples when implementing Dual-NDA. Generator SFID $\downarrow$ NIQE $\downarrow$ Diversity $\uparrow$ Label Score $\downarrow$ None 1.334 (0.531) 1.784 (0.065) 1.234 (0.209) 14.807 (14.297) ADCGAN 1.212 (0.481) 1.832 (0.055) 1.222 (0.180) 10.884 (11.041) ReACGAN 1.203 (0.499) 1.800 (0.078) 1.273 (0.171) 11.486 (10.904) CcGAN (ours) 1.114 (0.503) 1.738 (0.055) 1.251 (0.172) 11.809 (11.694) Ablation Study 2: Comparison of different combinations of NDA, Type I, and Type II negative samples. The best result for each metric is marked in boldface. Method SFID $\downarrow$ NIQE $\downarrow$ Diversity $\uparrow$ Label Score $\downarrow$ CcGAN (T-PAMI’23) 1.334 (0.531) 1.784 (0.065) 1.234 (0.209) 14.807 (14.297) w/ NDA (ICLR’21) 1.381 (0.527) 1.994 (0.081) 1.231 (0.167) 10.717 (10.371) w/ NDA+Type I 1.204 (0.564) 1.892 (0.063) 1.208 (0.179) 9.722 (9.897) w/ NDA+Type II 1.246 (0.545) 1.741 (0.045) 1.284 (0.198) 10.728 (9.789) w/ NDA+Type I & II 1.318 (0.570) 1.746 (0.058) 1.191 (0.226) 9.744 (9.227) w/ Type I 1.123 (0.478) 1.824 (0.045) 1.228 (0.206) 10.221 (10.301) w/ Type II 1.234 (0.598) 1.729 (0.074) 1.291 (0.186) 12.344 (12.153) w/ Type I & II (ours) 1.114 (0.503) 1.738 (0.055) 1.251 (0.172) 11.809 (11.694)
C.2 Example Type II Negative Samples
In this section, we present a selection of example Type II negative samples generated for both the UTKFace (128$\times$128) and Steering Angle (128$\times$128) experiments. The visualization of these samples ranges from Fig. 10 to Fig. 13. These visuals provide a clear illustration of the generated negative images. It is worth noting that a significant portion of these negative images exhibit evident distortions, artifacts, or are affected by overexposure issues.
Fig. 10: Some example Type II negative images for “Age=1” in the UTKFace (128$\times$128) experiment.
Fig. 11: Some example Type II negative images for “Age=60” in the UTKFace (128$\times$128) experiment.
Fig. 12: Some example Type II negative images with labels near $-40^{\circ}$ in the Steering Angle (128$\times$128) experiment.
Fig. 13: Some example Type II negative images with labels near $80^{\circ}$ in the Steering Angle (128$\times$128) experiment.
C.3 Example Fake Images from Baseline Methods
In this section, we showcase a selection of example synthetic images generated by both the baseline CcGAN and the proposed Dual-NDA methods. These examples are visualized in Fig. 14 for the UTKFace dataset (128$\times$128) and Fig. 15 for the Steering Angle dataset (128$\times$128).
@@ C.3 Example Fake Images from Baseline Methods
Fig. 14: Some example fake images from the baseline CcGAN and Dual-NDA in the UTKFace (128$\times$128) experiment.
Fig. 15: Some example fake images from the baseline CcGAN and Dual-NDA in the Steering Angle (128$\times$128) experiment.
C.4 Sampling Time Comparison
We also conduct a comparison of the sampling times for the baseline methods on the Steering Angle (64$\times$64) dataset, as depicted in Fig. 16. The visualization in the figure reveals that GAN-based methods require only a few seconds to generate 100,000 synthetic images, whereas diffusion models necessitate approximately 6 to 8 hours for the same task.
Fig. 16: Sampling time comparison on Steering Angle (64$\times$64). Each method is asked to generate 50 fake images for each of the 2000 distinct evaluation angles.

### 31 Dec 2023 v3

@@ C.2 Some extra ablation studies
Ablation Study 1: The effect of various generators for generating Type II
negative samples when implementing Dual-NDA.
Generator SFID $\downarrow$ NIQE $\downarrow$ Diversity $\uparrow$ Label Score $\downarrow$
None 1.334 (0.531) 1.784 (0.065) 1.234 (0.209) 14.807 (14.297)
ADCGAN 1.212 (0.481) 1.832 (0.055) 1.222 (0.180) 10.884 (11.041)
ReACGAN 1.203 (0.499) 1.800 (0.078) 1.273 (0.171) 11.486 (10.904)
CcGAN (ours) 1.114 (0.503) 1.738 (0.055) 1.251 (0.172) 11.809 (11.694)
Ablation Study 2: Comparison of different combinations of NDA, Type I, and Type II
negative samples. The best result for each metric is marked in boldface.
Method SFID $\downarrow$ NIQE $\downarrow$ Diversity $\uparrow$ Label Score $\downarrow$
CcGAN (T-PAMI’23) 1.334 (0.531) 1.784 (0.065) 1.234 (0.209) 14.807 (14.297)
w/ NDA (ICLR’21) 1.381 (0.527) 1.994 (0.081) 1.231 (0.167) 10.717 (10.371)
w/ NDA+Type I 1.204 (0.564) 1.892 (0.063) 1.208 (0.179) 9.722 (9.897)
w/ NDA+Type II 1.246 (0.545) 1.741 (0.045) 1.284 (0.198) 10.728 (9.789)
w/ NDA+Type I & II 1.318 (0.570) 1.746 (0.058) 1.191 (0.226) 9.744 (9.227)
w/ Type I 1.123 (0.478) 1.824 (0.045) 1.228 (0.206) 10.221 (10.301)
w/ Type II 1.234 (0.598) 1.729 (0.074) 1.291 (0.186) 12.344 (12.153)
w/ Type I & II (ours) 1.114 (0.503) 1.738 (0.055) 1.251 (0.172) 11.809 (11.694)
@@
C.3 Example Type II Negative Samples
In this section, we present a selection of example Type II negative samples generated for both the UTKFace (128$\times$128) and Steering Angle (128$\times$128) experiments. The visualization of these samples ranges from Figure 10 to Figure 13. These visuals provide a clear illustration of the generated negative images. It is worth noting that a significant portion of these negative images exhibit evident distortions, artifacts, or are affected by overexposure issues.
Figure 10: Some example Type II negative images for “Age=1” in the UTKFace (128$\times$128) experiment.
Figure 11: Some example Type II negative images for “Age=60” in the UTKFace (128$\times$128) experiment.
Figure 12: Some example Type II negative images with labels near $-40^{\circ}$ in the Steering Angle (128$\times$128) experiment.
Figure 13: Some example Type II negative images with labels near $80^{\circ}$ in the Steering Angle (128$\times$128) experiment.
C.4 Example Fake Images from Baseline Methods
In this section, we showcase a selection of example synthetic images generated by both the baseline CcGAN and the proposed Dual-NDA methods. These examples are visualized in Figure 14 for the UTKFace dataset (128$\times$128) and Figure 15 for the Steering Angle dataset (128$\times$128).
@@ C.4 Example Fake Images from Baseline Methods
Figure 14: Some example fake images from the baseline CcGAN and Dual-NDA in the UTKFace (128$\times$128) experiment.
Figure 15: Some example fake images from the baseline CcGAN and Dual-NDA in the Steering Angle (128$\times$128) experiment.
C.5 Sampling Time Comparison
We also conduct a comparison of the sampling times for the baseline methods on the Steering Angle (64$\times$64) dataset, as depicted in Figure 16. The visualization in the figure reveals that GAN-based methods require only a few seconds to generate 100,000 synthetic images, whereas diffusion models necessitate approximately 6 to 8 hours for the same task.
Figure 16: Sampling time comparison on Steering Angle (64$\times$64). Each method is asked to generate 50 fake images for each of the 2000 distinct evaluation angles.